Find the amount to which $500 will grow under each of the following conditions. Do not round intermediate calculations. Round your answers to the nearest cent.

6% compounded annually for 5 years. $

6% compounded semiannually for 5 years. $

6% compounded quarterly for 5 years. $

6% compounded monthly for 5 years.
Answer #1

Future value = Present value X (1+i=n)^nt Future value = $669.11 $500 X (1+6%/1)^(105) Future value = $671.96 $500 X (1+6%/2)